OpenStudy (anonymous):

What is the description of y=square root of 9x-3 to make it easy to graph using transformations of its parent function? A - the graph of y=3 square root of x, shifted right 3 units B - the graph of y=3 square root of x, shifted right 1/3 unit C- the graph of y= square root of x, shifted right 3 units and up 9 units D- the graph of y= square root of x, shifted right 1/3 units and up 9 units

OpenStudy (callisto):

\[y=\sqrt{9x-3}\]or\[y=\sqrt{9x}-3\]

OpenStudy (anonymous):

which answer choice?

OpenStudy (anonymous):

this question and its format confuses me

OpenStudy (callisto):

the graph mentioned in the question is the upper one or the lower one?

OpenStudy (anonymous):

it doesn't give me a graph

OpenStudy (callisto):

i meant y=square root of 9x-3

OpenStudy (anonymous):

oh its \[y=\sqrt{9x+3}\]

OpenStudy (anonymous):

-3*

OpenStudy (anonymous):

not + sorry

OpenStudy (anonymous):

@Callisto \[y=\sqrt{9x-3}\]

OpenStudy (campbell_st):

its B a zero exists when 9x - 3=0 or x = 1/3 which means the curve starts at x= 1/3 the comparison graph 3square root x has a zero at the origin